the measure of one of two adjacent angles is 116 degree more than thrice the other . find the measure of both angles if this sum is 140 degree

Answer by stanbon(75887)

the measure of one of two adjacent angles is 116 degree more than thrice the other . find the measure of both angles if this sum is 140 degree
------
Equation:
a + b = 140 degrees
a = 3b+116
-----
Substitute for "a" and solve for "b":
3b+116 + b = 140
4b = 24
b = 6 degrees
----
a + b = 140
a + 6 = 140
a = 134 degrees
==================
Cheers,
Stan H.
====================